Eve shift, 26 Jun 1999

Julie Whitmore (jaws@fnpxsr)
Mon, 28 Jun 1999 12:22:08 CDT

Sorry this is late, but I had a typo in the shiftsum email address. Here's
the shift summary for Sat evening. -julie

_____________________________________________________________________________

Shift Summary for Eve shift, 06/26/99

On Shift: Andrea Soddu, Julie

Hours of beam delivered at all intensities: 3.5hrs (2.5 w/ beam stop in)
Hours of beam delivered at nominal intensity: 1hr
Hours of E832 data (special runs not included): 1 hr
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: Problems with TeV injection, steering, losses, etc.

========================================================================
Runs ended this shift:
Run 13743 - BEAM_832 - beam debugging (TeV problems) - data no good
Run 13744 - BEAM_832 - good run (still going)

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift: Tevatron problems. Put beam stop in while MCR fixed
extraction problems and tuned beam.

========================================================================
Description of Shift:

run spill time comment
- ------------------------------------------------------------------------
13743 17:37 Start Beam 832 run.

17:46 1 spill of beam, then beam disappears. Called MCR.
There is a problem in a set of controllers near CDF
that is screwing up TeV injection. No estimate for beam.

126 20:28 Beam returns but is far off target.

130 20:32 Mary from MCR calls. Says they are having problems
controlling beam positions. We decide to close our
beam stop until they figure it out.

21:00 CA_out_of_limits alarm goes off because beam stop
is in.

22:10 Called MCR. Still having problems with losses in the
TeVatron. After they clear that up, they will work on
getting the beam centered on target. Hope to get us
beam w/i the hour.

23:00 MCR calls to say that the beam tune is better and we can
try opening our beam stop.

23:01 End run 13743.

13744 23:07 Begin run 13744. Noticed that L1 rates were too low
for NM2SEM value. Called MCR. They are aware of the
problem and are working on it. Hand off E832 run
to owl shift.

========================================================================
Items to Follow up on Next Shift

Continue taking BEAM_832 data. Do not take muon run or switch polarity of
NM4AN. Keep an eye on beam position on NM2WC1 and SEED2.